# Numerically Robust Fixed-Point Smoothing Without State Augmentation

### v1

@@ 1 Introduction
Bayesian filtering and smoothing enjoy numerous applications in areas like tracking, navigation, or control [Grewal and Andrews, 2014, Särkkä and Svensson, 2023]. They also serve as the computational backbone of popular machine learning methods, including message passing, sequence modelling, Gaussian processes, and probabilistic numerics [Grewal and Andrews, 2014, Särkkä and Solin, 2019, Murphy, 2023, Hennig et al., 2022]. Recent developments in Bayesian smoothing focus on subjects like linearisation [e.g. García-Fernández et al., 2016], temporal parallelisation [e.g. Särkkä and García-Fernández, 2020], or numerically robust implementations [e.g. Yaghoobi et al., 2022]. However, they all exclusively focus on fixed-interval smoothing, which targets the full time-series $p(x_{0:K}\mid y_{1:K})$ in $\mathcal{O}(K)$ memory, never on fixed-point smoothing, which only yields the initial value $p(x_{0}\mid y_{1:K})$ but does so in $\mathcal{O}(1)$ memory. Even though fixed-point smoothing is a crucial component of, for example, estimating past locations of a spacecraft [Meditch, 1969], it has yet to receive much attention in the literature on state estimation in dynamical systems. The experiments in Section 4 demonstrate fixed-point smoothing applications in probabilistic numerics and in a tracking problem.
@@
To close this gap, we introduce an implementation of numerically robust fixed-point smoothing. Our approach avoids the typical construction of fixed-point smoothers via state-augmentation [Biswas and Mahalanabis, 1972, Smith and Roberts, 1982], which increases the dimension of the state-space model and thus makes estimation needlessly expensive. And unlike previous work on fixed-point smoothing [Meditch, 1967a, Sarkka and Hartikainen, 2010, Rauch, 1963, Meditch, 1967b, 1976, Meditch and Hostetter, 1973, Nishimura, 1969], our perspective is not tied to any particular parametrisation of Gaussian variables. Instead, our proposed algorithm enjoys numerical robustness and compatibility with data streams, while maintaining minimal complexity. There exist tools for each of those desiderata, but our algorithm is the first to deliver them all at once. We demonstrate the algorithm’s efficiency and robustness on a sequence of test problems, including a probabilistic numerical method for boundary value problems [Krämer and Hennig, 2021], and show how to use fixed-point smoothing for parameter estimation in Gaussian state-space models.
@@ 2.1 Background on filtering and smoothing
This work only discusses linear, discrete state-space models with additive Gaussian noise because such models are the starting point for Bayesian filtering and smoothing. Nonlinear extensions are future work. For integers $d$ and $D$, let $A_{1},...,A_{K}\in\mathbb{R}^{D\times D}$ and $H_{1},...,H_{K}\in\mathbb{R}^{d\times D}$ be linear operators and $C_{0\mid 0},Q_{1},...,Q_{K}\in\mathbb{R}^{D\times D}$ and $R_{1},...,R_{K}\in\mathbb{R}^{d\times d}$ be covariance matrices. Introduce vectors $m_{0\mid 0},b_{1},...,b_{K}\in\mathbb{R}^{D}$ and $r_{1},...,r_{K}\in\mathbb{R}^{d}$. Assume that observations $y_{1:K}$ are available according to the state-space model (and potentially as a data stream)
@@
For example, the Kalman filter [Kalman, 1960] computes $p(x_{K}\mid y_{1:K})$ by initialising $p(x_{0}\mid y_{1:0})=p(x_{0})$ and alternating
@@ 2.2 Existing approaches to fixed-point smoothing
The state of the art in fixed-point smoothing does not use fixed-interval smoothers to compute $p(x_{0}\mid x_{1:K})$. There are two more efficient approaches: Either we solve the fixed-point smoothing problem by computing the solution to a specific high-dimensional filtering problem [Biswas and Mahalanabis, 1972], or we derive recursions for how $p(x_{0}\mid y_{1:k})$ evolves with $k=1,...,K$ [Meditch, 1967b, a, 1969]. Both solutions have advantages and disadvantages. We discuss both before explaining how our algorithm fixes their shortcomings.
@@
Suppose we temporarily ignore Cholesky-based forms and commit to covariance-based parametrisations of Gaussian distributions. In that case, we can improve the efficiency of the fixed-point smoother. It turns out that during the forward pass, the mean and covariance of $p(x_{0}\mid y_{1:k})$ follow a recursion that can be run parallel to the Kalman filter pass [Meditch, 1969, Sarkka and Hartikainen, 2010, Meditch, 1967b]. Sarkka and Hartikainen [2010] explain how to implement this recursion: start by initialising $m_{0\mid 0}$ and $C_{0\mid 0}$ according to Equation 1 and set $G_{0\mid 0}=I$. Then, iterate from $(m_{0\mid k-1},C_{0\mid k-1},G_{0\mid k-1})$ to $(m_{0\mid k},C_{0\mid k},G_{0\mid k})$ using the recursion [Sarkka and Hartikainen, 2010, Särkkä, 2013]
@@
Next to $(m_{0\mid k-1},C_{0\mid k-1},G_{0\mid k-1})$, these formulas only depend on the output of the prediction step as well as the smoothing gains $G_{k-1\mid k}$ (smoothing gains: Appendix B). The recursion in Equation 7 can be implemented to run simultaneously with the forward filtering pass. Unfortunately, even though this implementation avoids doubling the size of the state-space model and enjoys $\mathcal{O}(D^{2})$ memory and $\mathcal{O}(KD^{3})$ runtime, it does not have a Cholesky-based formulation. Thus, it cannot be applied to problems where numerical robustness is critical, like the probabilistic numerical simulation of differential equations [Krämer, 2024]. The main contribution of this paper is to generalise Equation 7 in a way that enables Cholesky-based parametrisations:
@@ 3.1 A new recursion for fixed-point smoothing
    Marginalise $p(x_{0}\mid y_{1:K}$ from $p(x_{K}\mid y_{1:K})$ and $p(x_{0}\mid x_{K},y_{1:K-1})$ via Equation 8.
@@ 3.2 Cholesky-based implementation
The missing link for Algorithm 1 is the merging of two affinely related conditionals in Equation 12. Since both conditionals in Equation 12 are affine and Gaussian, their combination $p(x_{0}\mid x_{k},y_{1:k})$ is Gaussian and available in closed form for all $k=1,...,K$. Its parametrisation depends on the parametrisation of each input distribution. For covariance- and Cholesky-based arithmetic, it looks as follows:
@@
Recall the initialisation of Algorithm 1 and the convention $y_{1:-1}=y_{1:0}=\emptyset$. For any $k=1,...,N$, if the fixed-point and smoothing conditionals are
@@
For any $k=1,...,N$ (and interpreting $y_{1:-1}=y_{1:0}=\emptyset$), if the previous fixed-point conditional and smoothing transition are
@@
$p_{k-1\mid k}\coloneqq$ $\,G_{0\mid k-1}p_{k-1\mid k}+p_{0\mid k-1},$ (16c)
$L_{P_{k-1\mid k}}\coloneqq$ $\,\mathfrak{R}^{\top},$ (16d)
@@ 4.1 Experiment I: How efficient is the fixed-point recursion?
We choose $K=1,000$, vary $d$, set the size of the hidden state to $D=2d$, and use Cholesky-based arithmetic for all estimators. We take Equation 1 and introduce a nonzero bias in all noises in Equation 2. Then, we randomly populate all system matrices in the state-space model with independent samples from $\mathcal{N}(0,\nicefrac{{1}}{{K^{2}}})$. Afterwards, we sample $y_{1:K}$ from this state-space model to generate toy data. A $K$-dependent covariance controls that the samples $y_{1:K}$ remain finite in 32-bit floating-point arithmetic. However, the precise values of the model and data do not matter for this experiment – only their size does. Finally, we vary $d$ and measure the runtime and memory requirements of each of the three methods. To measure runtime, we time the execution of the fixed-point smoothers: wall time in seconds, choosing the best of three runs. To measure memory consumption, we count the number of floating-point values the estimators carry from step to step, multiply this by 32 (because we use 32-bit arithmetic), and translate bits into bytes.
@@
Table 3: Runtime in seconds (wall-time, best of three runs). Lower is better. The column-wise lowest entries are bold and shaded. Randomly populated model. $K=1,000$ steps.
@@
Table 4: Memory in bytes (we use 32-bit arithmetic). Lower is better. The column-wise lowest entries are bold and shaded. Randomly populated model. $K=1,000$ steps.
@@ 5 Discussion
Our new approach to fixed-point smoothing strictly generalises existing techniques, but inherits some of their limitations: Even though the new recursion in Algorithm 1 is independent of the type of state-space model, implementing the fixed-point smoother in closed form assumes a linear Gaussian setup. Future work should explore robust fixed-point smoothing in nonlinear state-space models, for example, through posterior linearisation [García-Fernández et al., 2016]. Like all Gaussian smoothing algorithms, our methods have cubic complexity in the state-space dimension because of the matrix-matrix arithmetic or QR decompositions, respectively. Finally, while other parametrisations of Gaussian variables may be used instead of covariance or Cholesky-based parametrisations, such directions are future work.
@@ Acknowledgements
This work was supported by a research grant (42062) from VILLUM FONDEN. The work was partly funded by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606).

### v2

@@ 1 Introduction
Linear Gaussian state-space models and Bayesian filtering and smoothing enjoy numerous applications in areas like tracking, navigation, or control [Grewal and Andrews, 2014, Särkkä and Svensson, 2023]. They also serve as the computational backbone of contemporary machine learning methods that revolve around time-series, streaming data, or sequence modelling; for example, message passing, Gaussian processes, and probabilistic numerics [Grewal and Andrews, 2014, Särkkä and Solin, 2019, Murphy, 2023, Hennig et al., 2022]. Filtering and smoothing have also been used for constructing and training neural networks [Singhal and Wu, 1988, Gu and Dao, 2023, Chang et al., 2023] or continual learning [Sliwa et al., 2024]. All of these applications require fast and robust algorithms for state-space models; we propose a new one in this paper.
Recent developments in Bayesian smoothing focus on subjects like linearisation [e.g. García-Fernández et al., 2016], temporal parallelisation [e.g. Särkkä and García-Fernández, 2020], or numerically robust implementations [e.g. Yaghoobi et al., 2022]. However, they all exclusively focus on fixed-interval smoothing, which targets the full time-series $p(x_{0:K}\mid y_{1:K})$ in $\mathcal{O}(K)$ memory, never on fixed-point smoothing, which only yields the initial value $p(x_{0}\mid y_{1:K})$ but does so in $\mathcal{O}(1)$ memory. Even though fixed-point smoothing is a crucial component of, for example, estimating past locations of a spacecraft [Meditch, 1969], it has yet to receive much attention in the literature on state estimation in dynamical systems. The experiments in Section 4 demonstrate fixed-point smoothing applications in probabilistic numerics and in a tracking problem. We anticipate that probabilistic numerical solvers for differential equations will especially benefit from improving the practicality of fixed-point smoothers. The reason is that these algorithms closely connect to filtering and smoothing algorithms [Schober et al., 2019], especially the numerically robust kind [Krämer and Hennig, 2024], and that unknown initial conditions are typical for parameter estimation problems in differential equations and scientific machine learning [Rackauckas et al., 2020]. Section 4.2 revisits probabilistic numerics as a prime application for fixed-point smoothing.
@@
We introduce an implementation of numerically robust fixed-point smoothing. Our approach avoids the typical construction of fixed-point smoothers via state-augmentation [Biswas and Mahalanabis, 1972, Smith and Roberts, 1982], which increases the dimension of the state-space model and thus makes estimation needlessly expensive. And unlike previous work on fixed-point smoothing [Meditch, 1967a, Särkkä and Hartikainen, 2010, Rauch, 1963, Meditch, 1967b, 1976, Meditch and Hostetter, 1973, Nishimura, 1969], our perspective is not tied to any particular parametrisation of Gaussian variables. Instead, our proposed algorithm enjoys numerical robustness and compatibility with data streams, while maintaining minimal complexity. There exist tools for each of those desiderata, but our algorithm is the first to deliver them all at once. We demonstrate the algorithm’s efficiency and robustness on a sequence of test problems, including a probabilistic numerical method for boundary value problems [Krämer and Hennig, 2021], and show how to use fixed-point smoothing for parameter estimation in Gaussian state-space models.
@@ 2.1 Background on filtering and smoothing
This work only discusses linear, discrete state-space models with additive Gaussian noise because such models are the starting point for Bayesian filtering and smoothing. Nonlinear extensions are future work. For integers $d$ and $D$, let $A_{1},...,A_{K}\in\mathbb{R}^{D\times D}$ and $H_{1},...,H_{K}\in\mathbb{R}^{d\times D}$ be linear operators and $C_{0\mid 0},B_{1},...,B_{K}\in\mathbb{R}^{D\times D}$ and $R_{1},...,R_{K}\in\mathbb{R}^{d\times d}$ be covariance matrices. Introduce a vector $m_{0\mid 0}\in\mathbb{R}^{D}$. Assume that observations $y_{1:K}$ are available according to the state-space model (and potentially as a data stream)
@@
For example, the Kalman filter [Kalman, 1960] computes $\{p(x_{k}\mid y_{1:k})\}_{k=1}^{K}$ by initialising $p(x_{0}\mid y_{1:0})=p(x_{0})$ and alternating
@@ 2.2 Existing approaches to fixed-point smoothing
The state of the art in fixed-point smoothing does not use fixed-interval smoothers to compute $p(x_{0}\mid x_{1:K})$. There are two more efficient approaches: Either we solve the fixed-point smoothing problem by computing the solution to a specific high-dimensional filtering problem [Biswas and Mahalanabis, 1972, “fixed-point smoothing via state-augmented filtering” below], or we derive recursions for how $p(x_{0}\mid y_{1:k})$ evolves with $k=1,...,K$ [Meditch, 1967b, a, 1969, “fixed-point recursions” below]. Both solutions have advantages and disadvantages. We discuss both before explaining how our algorithm fixes their shortcomings.
@@
Suppose we temporarily ignore Cholesky-based forms and commit to covariance-based parametrisations of Gaussian distributions. In that case, we can improve the efficiency of the fixed-point smoother. It turns out that during the forward pass, the mean and covariance of $p(x_{0}\mid y_{1:k})$ follow a recursion that can be run parallel to the Kalman filter pass [Meditch, 1969, Särkkä and Hartikainen, 2010, Meditch, 1967b]. Särkkä and Hartikainen [2010] explain how to implement this recursion: start by initialising $m_{0\mid 0}$ and $C_{0\mid 0}$ according to Equation 1 and set $G_{0\mid 0}=I$. Then, iterate from $(m_{0\mid k-1},C_{0\mid k-1},G_{0\mid k-1})$ to $(m_{0\mid k},C_{0\mid k},G_{0\mid k})$ using the recursion [Särkkä and Hartikainen, 2010, Särkkä, 2013]
@@
Next to $(m_{0\mid k-1},C_{0\mid k-1},G_{0\mid k-1})$, these formulas only depend on the output of the prediction step as well as the smoothing gains $G_{k-1\mid k}$ (smoothing gains: Appendix B). For $k=1$, the fixed-point gain intentionally equals the smoothing gain, which Equation 7 expresses by initialising $G_{0\mid 0}=I$. The recursion in Equation 7 can be implemented to run simultaneously with the forward filtering pass. Unfortunately, even though this implementation avoids doubling the size of the state-space model and enjoys $\mathcal{O}(D^{2})$ memory and $\mathcal{O}(KD^{3})$ runtime, a Cholesky-based formulation has been unknown (until now). Thus, Equation 7 cannot be applied to problems where numerical robustness is critical, like the probabilistic numerical simulation of differential equations [Krämer, 2024]. The main contribution of this paper is to generalise Equation 7 to enable Cholesky-based parametrisations:
@@ 3.1 A new recursion for fixed-point smoothing
    Marginalise $p(x_{0}\mid y_{1:K})$ from $p(x_{K}\mid y_{1:K})$ and $p(x_{0}\mid x_{K},y_{1:K-1})$ via Equation 8.
@@ 3.2 Cholesky-based implementation
The missing link for Algorithm 1 is the merging of two affinely related conditionals in Equation 12. Since both conditionals in Equation 12 are affine and Gaussian, their combination $p(x_{0}\mid x_{k},y_{1:k-1})$ is Gaussian and available in closed form for all $k=1,...,K$. Its parametrisation depends on the parametrisation of each input distribution. For covariance- and Cholesky-based arithmetic, it looks as follows:
@@
Recall the initialisation of Algorithm 1 and the convention $y_{1:-1}=y_{1:0}=\emptyset$. For any $k=1,...,K$, if the fixed-point and smoothing conditionals (see Equation 4 & Appendix B for the latter) are parametrised by
@@
For any $k=1,...,K$ (again $y_{1:-1}=y_{1:0}=\emptyset$), if the previous fixed-point conditional and smoothing transition (see Equation 4 & Appendix B) are
@@
$p_{0\mid k}\coloneqq$ $\,G_{0\mid k-1}p_{k-1\mid k}+p_{0\mid k-1},$ (16c)
$L_{P_{0\mid k}}\coloneqq$ $\,\mathfrak{R}^{\top},$ (16d)
@@ 4.1 Experiment I: How efficient is the fixed-point recursion?
We choose $K=1,000$, vary $d$, set the size of the hidden state to $D=2d$, and use Cholesky-based arithmetic for all estimators. We take Equation 1 and introduce a nonzero bias in all noises in Equation 2. Then, we randomly populate all system matrices in the state-space model with independent samples from $\mathcal{N}(0,\nicefrac{{1}}{{K^{2}}})$. Afterwards, we sample $y_{1:K}$ from this state-space model to generate toy data. A $K$-dependent covariance controls that the samples $y_{1:K}$ remain finite in 32-bit floating-point arithmetic. However, the precise values of the model and data do not matter for this experiment – only their size does. Finally, we vary $d$ and measure the runtime and memory requirements of each of the three methods. To measure runtime, we use wall time in seconds. We display the best of three runs instead of the average because all codes are deterministic (no data-dependence, no randomness), thus the main driver for runtime differences is background machine noise. Selecting the best of three runs minimises this noise as much as possible. To measure memory consumption, we count the number of floating-point values the estimators carry from step to step, multiply this by 32 (because we use 32-bit arithmetic), and translate bits into bytes.
@@
Table 3: Runtime in seconds (wall-time, best of three runs). Lower is better. The column-wise lowest are bold & shaded. Randomly populated model. $K=1,000$ steps. All methods use Cholesky-based parametrisations.
@@
Table 4: Memory in bytes (we use 32-bit arithmetic). Lower is better. The column-wise lowest entries are bold and shaded. Randomly populated model. $K=1,000$ steps. All methods use Cholesky-based parametrisations.
@@ 5 Discussion
Our new approach to fixed-point smoothing strictly generalises existing techniques, but inherits some of their limitations: Even though the new recursion in Algorithm 1 is independent of the type of state-space model, implementing the fixed-point smoother in closed form assumes a linear Gaussian setup. Future work should explore robust fixed-point smoothing in nonlinear state-space models, for example, through posterior linearisation [García-Fernández et al., 2016]. Like all Gaussian smoothing algorithms, our methods have cubic complexity in the state-space dimension because of the matrix-matrix arithmetic or QR decompositions, respectively. Finally, other parametrisations of Gaussian variables could be used with Algorithm 1 instead of covariance or Cholesky-based parametrisations (Algorithms 2 and 3); for example, the information form or ensembles [Murphy, 2022, Houtekamer and Mitchell, 2005]. Combining these alternatives with numerically robust fixed-point smoothing could be interesting avenues for future work.
@@ Acknowledgements
This work was supported by a research grant (42062) from VILLUM FONDEN. The work was partly funded by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606). This project received funding from the European Research Council (ERC) under the European Union’s Horizon programme (grant agreement 101125993).